Session: Sep 1999 (session 3)
Wavelength: 30 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (A. Kraus at Fri Sep 10 08:44:11 METDST 1999) Success

Torun (K.Borkowski at Fri Sep 10 10:19:46 METDST 1999) Failure
Not observed due to ongoing repair of the terminal.

Westerbork (Tony Foley at Fri Sep 10 18:03:42 METDST 1999) Success
Array re-phased shortly before start of run. Used 12 telescopes

Onsala (A. Polatidis at Sat Sep 11 13:08:28 METDST 1999) Success

Jodrell (Alastair Gunn at Wed Sep 15 12:13:22 METDST 1999) Minor failures
Telescope reached cable wrap at approximately 08:05 UT. Some data on 3C84 were lost due to unwrapping. Back on source at 08:18 UT. LO glitched at about 13:00 UT, reset after 20 minutes. Caltemp values in this experiment are wrong. No other known problems.

Calibration (Jean\-Francois Desmurs at Sun Nov 7 15:19:38 MET 1999) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: eo004b.antab. Read the comments in the file to get further information. (Rem : Jb no Tsys values)